OpenStudy (anonymous):

Is the simplified form of 2 Radical3 -2 radical3, RRATIONAL??? yes or no Problem #2 Is the simplified form of 2 radical3 * rational 12, RATIONAL????? yes or No

OpenStudy (kirbykirby):

\[2\sqrt{3}-2\sqrt{3}\] ?

OpenStudy (anonymous):

yes, can you help me with both problems please!!!

OpenStudy (kirbykirby):

sure. Well the first one you re subtracting off the same number itself, what does that give :) ?

OpenStudy (anonymous):

0

OpenStudy (kirbykirby):

yes. Do you know if 0 is rational?

OpenStudy (kirbykirby):

recall that a rational number is a number that can be expressed as a fraction of integers, \[ \frac{a}{b}\], where \(b \ne 0\)

OpenStudy (kirbykirby):

\(2\sqrt{3}\times \sqrt{12}=2\sqrt{3}\times \sqrt{4\cdot 3}=2\sqrt{3}\times 2\sqrt{3}\)